imbadkidjenna:

5x + 10 = 15 0 3 O 1 0 5 O 25

darkknight:

5x + 10 = 15 5x =5 x = ?

Convert:

5x + 10 = 15 -10 -10 5x = 5 5x/5 = 5/5 x = 1

LimitedEditionShawty:

First, you Subtract 10 from both sides of the equation, then you simplify it then divide both sides of the equation by the same term, next you simplify again so x=1.

supie:

Yeah, \[5x+10=15\] -10 to both sides \[5x+10−10=15−10\]\[5x=5 \] ÷5 to both sides (fraction bar is division. \(\frac{5x}{5}=\frac{5}{5}\) \(x=?\)

ramen:

And btw convert, that is a direct answer.

smackz:

ok what i would do in this is first 5X + 10 = 15 - 10 = -10 5x = 5 x = 1 becuz u divide 5 by 5x which cancels out then 5 by 5 which comes to one hope this helps :D

Convert:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @ramen And btw convert, that is a direct answer. \(\color{#0cbb34}{\text{End of Quote}}\) How..? I explained how I reached that answer with the work shown in each step I made.

ramen:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Convert \(\color{#0cbb34}{\text{Originally Posted by}}\) @ramen And btw convert, that is a direct answer. \(\color{#0cbb34}{\text{End of Quote}}\) How..? I explained how I reached that answer with the work shown in each step I made. \(\color{#0cbb34}{\text{End of Quote}}\) Darkknight already did

darkknight:

no no one gave a direct answer, its fine. Why are there 5 of the same answers though? 2 is understandable but like...

Convert:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @ramen \(\color{#0cbb34}{\text{Originally by}}\) @Convert \(\color{#0cbb34}{\text{Originally Posted by}}\) @ramen And btw convert, that is a direct answer. \(\color{#0cbb34}{\text{End of Quote}}\) How..? I explained how I reached that answer with the work shown in each step I made. \(\color{#0cbb34}{\text{End of Quote}}\) Darkknight already did \(\color{#0cbb34}{\text{End of Quote}}\) My bad, I was already typing my reply and I guess he sent it a little before me.

ramen:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Convert \(\color{#0cbb34}{\text{Originally Posted by}}\) @ramen \(\color{#0cbb34}{\text{Originally by}}\) @Convert \(\color{#0cbb34}{\text{Originally Posted by}}\) @ramen And btw convert, that is a direct answer. \(\color{#0cbb34}{\text{End of Quote}}\) How..? I explained how I reached that answer with the work shown in each step I made. \(\color{#0cbb34}{\text{End of Quote}}\) Darkknight already did \(\color{#0cbb34}{\text{End of Quote}}\) My bad, I was already typing my reply and I guess he sent it a little before me. \(\color{#0cbb34}{\text{End of Quote}}\) It's alright, we got the answer ig

imbadkidjenna:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @darkknight well whatever, @imbadkidjenna got ur answer? \(\color{#0cbb34}{\text{End of Quote}}\) yea thanks

darkknight:

np